OpenStudy (anonymous):

7-3y = t^2 Differentiate w.r.t to y on both sides @ganeshie8 @Johnbc

ganeshie8 (ganeshie8):

\[\large \dfrac{d}{dy} \left(7-3y\right) = \dfrac{d}{dy}\left(t^2\right)\]

OpenStudy (aum):

\( \Large \frac {d}{dy}(7-3y) = \frac {d}{dy}t^2 \\ \Large -3 = 2t\frac {dt}{dy} \)

ganeshie8 (ganeshie8):

7 is a number constant, whats the derivative of a number with respect to variable ?

ganeshie8 (ganeshie8):

does the value of 7 change as the variable y changes ?

OpenStudy (anonymous):

I got that , but R.H.S

ganeshie8 (ganeshie8):

you have a different variable on RHS, so you need to use chain rule

OpenStudy (anonymous):

@Johnbc told we treat it as constant

OpenStudy (anonymous):

In integration.

ganeshie8 (ganeshie8):

we treat it as constant if t does not depend on y, but here we have `t = f(y) `

OpenStudy (anonymous):

You have 2 different variables multiplied by each other in this problem so for the derivative you would need to apply the chain rule

OpenStudy (anonymous):

ohk , thanks